OpenStudy (anonymous):

Solve using the elimination method 7x + 7y = 7 x + 2y = 2 pleas explain I do not get this at all

OpenStudy (akshay_budhkar):

ok let us go step by step

OpenStudy (anonymous):

step one is do divide the first equation by 7 to make your life simpler. you get \[x+y=1\] and now i shut up

OpenStudy (akshay_budhkar):

\[7x + 7y- 7 \therefore x+y =1 \therefore x=y-1\]

OpenStudy (precal):

7x+7y=7 x+2y=2 The goal is to eliminate the x's or the y's It is your choice

OpenStudy (precal):

or you could mutiply -7 to the second equation

OpenStudy (anonymous):

ok now i put in my two cents again. says "elimination" not "substitution" although i would use the second method as well

OpenStudy (precal):

Technically the method does not matter but since it did say elimination it is better to cancel out the x's or the y's I am choosing to get rid of the x's by multiplying -7 to equation 2

OpenStudy (akshay_budhkar):

substituting that in equation 2 we have\[\begin {array} x x+2y=2 \\ y-1+2y=2 \\ 3y=3 \\ y=1 \end {array}\]

OpenStudy (akshay_budhkar):

opps my bad

OpenStudy (akshay_budhkar):

didnt read elimination sorry

OpenStudy (akshay_budhkar):

ok in this you can solve by subtracting the two equations that is x+y=1 and x+2y=2.. you get 2y-y=2-1 therefore y=1

OpenStudy (anonymous):

ok I am so confused

OpenStudy (anonymous):

ok I get it sorry didnt get it at first

OpenStudy (akshay_budhkar):

so is it fine now?

OpenStudy (anonymous):

Yes it is thank you so much I get it now i hope

OpenStudy (akshay_budhkar):

if u have any doubt clear it =D all the best

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

I thought it would be 2y-y=2/1

OpenStudy (anonymous):

there for y=2

OpenStudy (precal):

I will do it for you using elimination

OpenStudy (anonymous):

i just need the process explained to me it just is not clicking im sorry

OpenStudy (precal):

7x+7y=7 x+2y=2 7x+7y=7 -7(x+2y=2) 7x+7y=7 -7x-14y=-14 add equations together -7y=-7 divide by -7 to both sides y=1 x+2y=2 x+2(1)=2 x+2=2 -2 both sides x=0 solution x=0 and y=1

OpenStudy (anonymous):

ok this makes since to me thank you

OpenStudy (anonymous):

so this would be and infanite solution correct because it would graph as the same line not two intersecting line?

OpenStudy (precal):

Another name for this method is called multiplication and addition method First you pick a number to multiply one of the equations then you add both equations together goal is to get rid of the x's or the y's that is why is is called elimination. If you wanted to get rid of the the y's instead you would have to multiply 2 to the first equation and then -7 to the second equation this will produce 14y and -14y and then you would have solve for x first. Many ways to eliminate x's or y's

OpenStudy (precal):

No this is one solution because we were able to find the point of intersection

OpenStudy (anonymous):

oh ok so what would a infanite solution look like?

OpenStudy (anonymous):

sorry to ask so many questions, but I really want to grasp this and I am having a really hard time

OpenStudy (precal):

infinite solutions is the same line given in multiple form example x+y=4 2x+2y=8 did you notice they are the same I only multiply 2 to the original equation

OpenStudy (anonymous):

yes I see

OpenStudy (precal):

No solution is 2 parallel lines, they never cross. look for same slopes y=2x+3 y=2x+6

OpenStudy (anonymous):

thank you so much for explaining this I understand it so much better now than i did from the book

OpenStudy (precal):

The number of solutions is determined by the number of times the equations cross: One solution means they cross in one place (X,Y) No solution means they do not cross (Parallel lines) Infinitely many solution means they cross in many places (because they are the same line)

OpenStudy (anonymous):

Thank you so much

OpenStudy (precal):

Anytime :)